OpenStudy (anonymous):

I need help writing an equation given a slope and a point. I have to explain in words the steps involved and show an example.

OpenStudy (lgbasallote):

you'll have to give the slope and the point

OpenStudy (anonymous):

well first i have to explain the steps on how to solve this type on problem.

OpenStudy (lgbasallote):

you use the point-slope formula \[y - y_1 = m(x-x_1)\] where: y1 = y coordinate m = slope x1 = x coordinate

OpenStudy (anonymous):

okay but could you elaborate and say like step 1....... step 2 ........ do you get what i'm saying?

OpenStudy (anonymous):

@lgbasallote